Find the inverse of the function.

f(x) = 3x - 2

I'm struggling in this topic
1
Expert's answer
2014-05-10T08:12:10-0400
f(x)=3x-2.
f(x)+2=3x ---> x=(f(x)+2)/3
Inverse function : f-1(x)=(x+2)/3.